# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Abstract
Understanding the behavior of quantum many-body systems under decoherence is essential for developing robust quantum technologies. Here, we examine the fate of weak ergodicity breaking in systems hosting quantum many-body scars when subject to local pure dephasing—an experimentally relevant form of environmental noise. Focusing on a large class of models with an approximate su(2)-structured scar subspace, we show that scarred eigenmodes of the Liouvillean exhibit a transition reminiscent of spontaneous $\mathbb{PT}$-symmetry breaking as the dephasing strength increases. Unlike previously studied non-Hermitian mechanisms, this transition arises from a distinct quantum jump effect. Remarkably, in platforms such as the XY spin ladder and PXP model of Rydberg atom arrays, the critical dephasing rate shows only weak dependence on the system size, revealing an unexpected robustness of scarred dynamics in noisy environments.
@@
Introduction—Recent advances in quantum simulations [1, 2, 3, 4, 5] have opened a window to studying thermalization in isolated quantum many-body systems [6, 7, 8, 9, 10]. Many such systems are now understood to host atypical non-thermalizing eigenstates known as quantum many-body scars (QMBSs) [11, 12, 13], which are ubiquitous in many physical platforms, including Rydberg atom arrays [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], the Heisenberg-type spin models [29, 30, 31, 32], ultracold atoms in tilted optical lattices [33, 34, 35, 36, 37, 38], superconducting circuits [39, 40], and many other [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. An important class of QMBSs are those associated with a so-called ‘restricted su(2) spectrum generating algebra’ [60, 61, 62, 63], whereby scarred eigenstates distribute with nearly equal energy spacings and often have large overlap with product states. The unique properties of this class of scars are of interest in quantum information processing [64, 65, 66], e.g., they have been used to generate Greenberger-Horne-Zeilinger states [67] and disorder-tunable entanglement [68].
The existing platforms for quantum simulation, however, are prone to decoherence caused by the surrounding environment. The decoherence – particularly pure dephasing – can significantly impact individual qubits [69], limiting their useful lifetime to around $10$ $\mu$s in current state-of-the-art superconducting platforms [70]. While dissipation was studied in kinetically-constrained models [71, 72] and time crystals [73, 74, 75, 76], its impact on the robustness of QMBS states remains poorly understood [77, 78, 79]. For example, while Ref. [77] recently demonstrated a construction for embedding QMBS into decoherence-free subspaces, a comprehensive understanding of realistic decoherence effects, such as pure dephasing, remains elusive. Previous studies of QMBSs to perturbations in closed systems [80, 81] suggest that they might quickly thermalize in the presence of noise.
Figure 1: (a) Schematic of a general open quantum system with approximate scars. The scar and thermal subsystems lose their quantum coherence to the bath with the dephasing rate $\gamma$, while interacting via the coupling $\hat{V}$. (b)-(d) The Liouvillean spectral transition in the SCL model, Eq. (Liouvillean Spectral Transition in Noisy Quantum Many-Body Scars), for three different dephasing strengths $\gamma\!=\!0$, $0.0005$, and $0.02$. Red and blue dots represent scar (S) eigenvalues, while yellow and gray ones represent thermal (T) eigenvalues, labeled according to their inversion symmetry quantum number, $p\!=\!\pm 1$. In panel (d), $(l,s)$ represents the $s$th eigenvalue in the $l$th layer [see text below Eq. (4)]. The scar eigenvalues in the $l^{\prime\prime}$th layer undergo LPTS breaking as $\gamma$ increases, while those in the $l^{\prime}$th layer do not [compare red dots in panel (c) with (d)]. The data in (b)-(d) is obtained by exact diagonalization for the choice of couplings $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$, system size $L\!=\!10$ with open boundary conditions.
In this paper, we consider a large class of models possessing a scar subspace with an su(2) spectrum generating algebra and coupled to the environment via local pure dephasing $\gamma$ on each site, see Fig. 1(a). In the presence of dissipation, it will be important to distinguish the cases where the spectrum generating algebra is exact versus only approximately obeyed within the scar subspace. The former is characteristic of analytical models of QMBSs [60, 61, 62], while the latter describes various experimental realizations [14, 37, 39, 68]. In contrast to exact scars, we find that approximate scars undergo a process resembling spontaneous Liouvillean $\mathbb{PT}$-symmetry (LPTS) breaking at a non-zero dephasing rate $\gamma_{\star}^{\mathrm{S}}$. Notably, we will argue that $\gamma_{\star}^{\mathrm{S}}$ for approximate scars depends only polynomially on system size, in stark contrast with chaotic systems where $\gamma_{\star}$ is inversely proportional to the square of the Liouvillean eigenvalue density [82]. This difference in scaling suggests that approximate scars can maintain LPTS and dynamical scar signatures at finite dephasing rates in much larger systems than naïvely expected. We identify the origin of this phenomenon in the quantum jumper effect beyond previously considered non-Hermitian mechanisms [83, 84]. We illustrate our conclusions using the experimentally relevant model of superconducing qubit ladders [39], while in Appendix we provide further evidence for the PXP model of Rydberg atom arrays [14].
Quantum Liouvillean spectra—The Hamiltonian of a QMBS system can be written as $\hat{H}\!=\!\hat{S}_{0}\bigoplus\hat{T}\!+\!\hat{V}$, where $\hat{S}_{0}$ and $\hat{T}$ denote the Hamiltonian of scar and thermal subsystems, respectively, while $\hat{V}$ represents the coupling between them [11], as illustrated in Fig. 1(a). Exact scar case corresponds to $\hat{V}\!=\!\hat{0}$, while approximate scars correspond to $\hat{V}\neq\hat{0}$. The eigenvalues and eigenstates of an isolated system (with the total Hamiltonian $\hat{H}$) are denoted by $E_{n}$ and $|E_{n}\rangle$, $n\!=\!1,2,\cdots,D$, with $D$ being the Hilbert space dimension. Among these, the scar eigenstates are denoted by $\big{|}E_{n_{\mathrm{S}}}\big{\rangle}$, while $\big{|}E_{n_{\mathrm{T}}}\big{\rangle}$ denotes the remaining thermal eigenstates. Notice that these eigenstates cannot be directly obtained from the subsystem Hamiltonians in the approximate scar case.
We now consider each site to have individual decoherence to a bath. Under the Markov approximation, such a system is generally described by a Liouville superoperator [85]. Here, we focus on pure dephasing, with the corresponding superoperator
$\begin{split}\mathcal{L}&=-i\left(\hat{H}\otimes\hat{I}-\hat{I}\otimes\hat{H}^{*}\right)+\gamma\mathcal{D},\\ \mathcal{D}&=\mathcal{D}^{\prime}-\sum_{j}\hat{I}_{j}\otimes\hat{I}_{j},\quad\quad\mathcal{D}^{\prime}=\sum_{j}\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}\end{split}$ (1)
where $\gamma$ is the pure dephasing rate and $\hat{\sigma}^{z}_{j}$ is a local jump operator on $j$th lattice site. The operator $|E_{n}\rangle\langle E_{m}|$ in $\mathcal{L}$ has a vector representation given by $|E_{n}\rangle\!\otimes\!|E_{m}\rangle$ [85]. The $k$th eigenvalue and eigenmode of the Liouville superoperator in Eq. (1) are denoted by $\lambda_{k}$ and $|\lambda_{k}\rangle\!\rangle$, respectively, with $k=1,2,\cdots,D^{2}$.
To understand the spectral properties of Liouvillean, we perform a transformation [82] $\mathcal{L}^{\prime}\!=\!\mathcal{L}+\gamma\big{(}\mathrm{Tr}(\mathcal{D})/\mathrm{Tr}(\mathcal{I})\big{)}\mathcal{I}$, with $\mathcal{I}\!=\!\hat{I}\!\otimes\!\hat{I}$, which makes the Liouvillean spectrum traceless, i.e., sets the average of $\mathrm{Re}(\lambda^{\prime}_{k})$ to zero. For an isolated system ($\gamma\!=\!0$), the spectrum is purely imaginary, $\mathrm{Im}(\lambda_{k})\!=\!E_{n}{-}E_{m}$, and $|\lambda_{k}\rangle\!\rangle\!=\!|E_{n}\rangle\!\otimes\!|E_{m}\rangle$. As we turn on dephasing, the eigenvalues on the imaginary axis move towards each other and collide, jumping into the complex plane [82]. This general picture of Liouvillean spectral evolution is illustrated in Fig. 1(b)-(d) for the example of the SCL model that we study in detail next.
@@
SCL model—The SCL model describes spin-$1/2$ XY chains in a Cruetz ladder configuration [86, 39]. The Hamiltonian, $\hat{H}_{\mathrm{SCL}}\!=\!\hat{H}_{\mathrm{h,1}}+\hat{H}_{\mathrm{h,2}}+\hat{H}_{\mathrm{v}}+\hat{H}_{\mathrm{x}}$, is a sum of Hamiltonians $\hat{H}_{\mathrm{h,\alpha}}$ for each of the two horizontal legs $\alpha\in\{0,1\}$ of the ladder, the vertical coupling between them, $\hat{H}_{\mathrm{v}}$, and the cross-coupling $\hat{H}_{\mathrm{x}}$ present in experiment [39]. Specifically,
@@
$\hat{H}_{\mathrm{v}}$ $=$ $J\sum_{j}\left(\hat{\sigma}^{+}_{j,\bar{\alpha}}\hat{\sigma}^{-}_{j,\alpha}+\mathrm{h.c.}\right),$ (2)
$\hat{H}_{\mathrm{x}}$ $=$ $\sum_{\alpha}\sum_{j}J_{\mathrm{x},j}\left(\hat{\sigma}^{+}_{j,\alpha}\hat{\sigma}^{-}_{j+1,\bar{\alpha}}+\mathrm{h.c.}\right),$
@@
where $\hat{\sigma}^{\pm}_{j,\alpha}$ is the standard raising (lowering) spin-$1/2$ operators for $j$th spin in layer $\alpha$, with $\bar{\alpha}$ denoting the layer opposite to $\alpha$. The ladder contains $L=2N$ spins in total. We assume open boundary conditions and set $J=1$ as the overall energy scale. Note that the cross-coupling $J_{\mathrm{x},j}$ can be site-dependent or even random.
When $J_{\mathrm{x},j}\!=\!0$, the SCL model hosts exact scar eigenstates $|E_{n_{\mathrm{S}}}\rangle$, which are simply the Dicke states of a free collective spin, $\hat{S}_{0}\!=\!\sum_{j}\hat{\sigma}^{x}_{j}$ [87]. These eigenstates overlap with the product state $\left|\Pi\right\rangle\!=\!\left|\begin{subarray}{c}\bullet\bullet\bullet\\ \circ\circ\circ\end{subarray}\cdots\right\rangle$, where $\bullet$” and $\circ$” denote the basis states of a single site. The state $|\Pi\rangle$ exhibits perfect revivals for $J_{\mathrm{x},j}\!=\!0$. On the other hand, when $J_{\mathrm{x},j}\!\neq\!0$, the QMBS eigenstates $\big{|}E_{n_{\mathrm{S}}}\big{\rangle}$ are only approximately given by the Dicke states and the revivals acquire a decaying envelope [39]. In the Liouvillean spectrum, QMBS eigenmodes can be identified via the overlap with QMBS eigenmodes $|E_{n_{\mathrm{S}}}\rangle\!\otimes\!|E_{m_{\mathrm{S}}}\rangle$ at $\gamma\!=\!0$. When the dephasing becomes strong, one can only track the scar states by following smooth changes of the Liouville spectrum as $\gamma$ is slowly varied.
Spectral transition in SCL model—In Fig. 1(b)-(d), we present examples of Liouvillean spectra of the SCL model in Eq. (Liouvillean Spectral Transition in Noisy Quantum Many-Body Scars) with approximate scars ($J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$) at dephasing-free, small, and large pure dephasing rates, respectively. The QMBSs display distinct real Liouvillean eigenvalues. At weak dephasing $\gamma\!=\!0.0005$, those normalized eigenvalues featuring substantial displacement [red dots in panel (c)] cluster on the imaginary axis. In contrast, at the higher dephasing rate of $\gamma\!=\!0.02$ [panel (d)], the same eigenvalues exhibit a qualitatively different distribution pattern in the form of a sparse diamond grid. We also note that the right-most eigenmodes at larger dephasing rate, Fig. 1(d), have large overlap with $|\rho_{\Pi}\rangle\!\rangle=|\Pi\rangle\!\otimes\!|\Pi\rangle$ [87].
The stark difference between panels (c) and (d) of Fig. 1 reveals a shift of the eigenvalues along the real axis, which can be understood from the symmetry classification developed in Ref. [88]. The SCL Liouvillean obeys the $\mathbb{PT}$ symmetry property, which is generated by the superoperator $\mathcal{T}_{-}$,
$\mathcal{T}_{-}=\sqrt{\pm 1}\Big{[}\Big{(}\prod_{i}\hat{\sigma}_{i}^{x}\Big{)}\otimes\hat{I}\Big{]}\mathcal{K},\quad\quad\mathcal{T}_{-}\mathcal{L}\mathcal{T}_{-}^{-1}=-\mathcal{L},$ (3)
where $\mathcal{K}$ is the complex-conjugation superoperator. The existence of $\mathcal{T}_{-}$ is responsible for $D_{2}$ dihedral symmetry of the QMBS Liouvillean spectrum in Fig. 1(c) [82]. At $\gamma\!=\!0$, $|-(\lambda^{\prime})^{\ast}\rangle\!\rangle$ degenerates with $|\lambda^{\prime}\rangle\!\rangle$ and these lie on the imaginary axis. Beyond a certain $\gamma$, $\mathbb{P}\mathbb{T}|\lambda^{\prime}\rangle\!\rangle\!=\!|\!-\!(\lambda^{\prime})^{\ast}\rangle\!\rangle$ continues to hold, but $\lambda^{\prime}$ and $-(\lambda^{\prime})^{\ast}$ are no longer the same. Thus, after this critical point, pairs of degenerate eigenvalues move symmetrically away from the imaginary axis along the real axis, resulting in LPTS breaking [82].
A few comments are in order. The LPTS breaking described above occurs only in inversion symmetry sector $p\!=\!\mp 1$, corresponding to odd or even $N$, respectively, see $N\!=\!5$ case in Fig. 1(c). In contrast, the LPTS is already broken for infinitesimal $\gamma$ in $p\!=\!\pm 1$ sector, also correlating with odd or even $N$. Moreover, the spectral features in Fig. 1 are exclusive to models with approximate scars, while exact scars break LPTS at infinitesimal $\gamma$. Intuitively, the latter is due to the exact energy spacing, which leads to the degeneracy of energy differences, $E_{n_{S}}{-}E_{m_{S}}$, on the imaginary axis. Hence, the Liouvillean eigenvalues of exact scars jump into the complex plane for any $\gamma\!\neq\!0$, as confirmed in the SM for $J_{\mathrm{x},j}\!=\!0$ [87].
@@
The process of LPTS breaking can be captured more quantitatively by sweeping the dephasing rate $\gamma$. Figures 2(a)-(b) illustrate the imaginary part of eigenvalues $\lambda$ and the real part of normalized (shifted) eigenvalues $\lambda^{\prime}/\gamma$ as a function of $\gamma$. At weak dephasing, several eigenvalues with inversion symmetry $p\!=\!-1$ remain stationary along the real axis, while moving along the imaginary one. As $\gamma$ increases, some pairs of eigenvalues collide on the imaginary axis and abruptly bifurcate into two branches at specific dephasing rates $\gamma_{\star}^{\mathrm{S(1,2,3)}}$ along the real axis. These critical rates can be accurately estimated in a perturbation theory framework discussed below, whose prediction is shown in Fig. 2(c).
LPTS breaking can also be seen in the average movement velocity $\mathrm{Re}(\bar{v})\!=\!\mathrm{d}\!\left\langle|\mathrm{Re}(\lambda^{\prime}/\gamma)|\right\rangle/\mathrm{d}\gamma$ as a function of $\gamma$ in Fig. 2(d). For the $p$=$-1$ sector, a few distinct peaks align with the fast breaking velocities in Figs. 2(a)-(b) and the breaking points $\gamma_{\star}^{\mathrm{S(1,2,3)}}$ correspond to the front of three peaks. There are no such abrupt peaks in $p$=$+1$ sector of QMBSs. In contrast, $\mathrm{Re}(\bar{v})$ of thermal eigenvalues becomes smooth with increasing $\gamma$.

### v3

@@ Abstract
Understanding the behavior of quantum many-body systems under decoherence is essential for developing robust quantum technologies. Here, we examine the fate of weak ergodicity breaking in systems hosting quantum many-body scars when subject to local pure dephasing—an experimentally relevant form of environmental noise. Focusing on a large class of models with an approximate su(2)-structured scar subspace, we show that scarred eigenmodes of the Liouvillian exhibit a transition reminiscent of spontaneous $\mathbb{PT}$-symmetry breaking as the dephasing strength increases. Unlike previously studied non-Hermitian mechanisms, this transition arises from a distinct quantum jump effect. Remarkably, in platforms such as the XY spin ladder and PXP model of Rydberg atom arrays, the critical dephasing rate shows only weak dependence on the system size, revealing an unexpected robustness of scarred dynamics in noisy environments.
@@
Introduction—Recent advances in quantum simulations [1, 2, 3, 4, 5] have opened a window to studying thermalization in isolated quantum many-body systems [6, 7, 8, 9, 10]. Many such systems are now understood to host atypical non-thermalizing eigenstates known as quantum many-body scars (QMBSs) [11, 12, 13]. QMBSs are ubiquitous in many physical systems, including Rydberg atom arrays [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], the Heisenberg-type spin models [29, 30, 31, 32], ultracold atoms in tilted optical lattices [33, 34, 35, 36, 37, 38], superconducting circuits [39, 40], and many others [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. The unique properties of QMBSs are also of interest in quantum information processing [60, 61, 62], e.g., they have been used to prepare Greenberger-Horne-Zeilinger states [63].
[…]
In this paper, we consider a large class of QMBS models coupled to the environment via local pure dephasing $\gamma$ on each site. We focus on models with a so-called restricted su(2) spectrum generating algebra [77, 78, 79, 80], which includes experimentally-realized QMBSs in several platforms [14, 37, 39, 81]. In the absence of dephasing, these QMBS eigenstates distribute with nearly equal energy spacings and the system exhibits long-lived coherent dynamics when prepared in special initial states. As the dephasing increases beyond a critical value $\gamma_{\star}$, the thermal bulk of the spectrum undergoes a spontaneous Liouvillian $\mathbb{PT}$-symmetry (LPTS) breaking transition [82]. This occurs rapidly, as the critical dephasing is inversely proportional to the square of the Liouvillian eigenvalue density [82] and hence exponentially small in system size, $\gamma_{\star}\sim\exp(-L)$. By contrast, we find that QMBS eigenvalues undergo a distinct spectral transition at $\gamma_{\star}^{\mathrm{S}}$, which scales polynomially with system size, $\gamma_{\star}^{\mathrm{S}}\sim L^{-1}$. This difference in scaling has a striking manifestation for the robustness of QMBS signatures under dissipative dynamics. We identify the origin of this phenomenon in the quantum jumper effect beyond previously considered non-Hermitian mechanisms [83, 84]. We illustrate our conclusions using an experimentally relevant model of superconducing qubits from Refs. [39, 81], while in End Matter we provide further evidence for the PXP model of Rydberg atom arrays [14].
@@
Creutz Ladder (CL) model—We consider a paradigmatic QMBS model describing superconducting qubits in a Creutz ladder configuration [85, 39]. A variant of this model was used to experimentally demonstrate QMBS-enabled tunability of entanglement in the presence of disorder [81]. The total Hamiltonian, $\hat{H}=\hat{H}_{\mathrm{CL}}\!=\!\hat{H}_{\mathrm{h,1}}+\hat{H}_{\mathrm{h,2}}+\hat{H}_{\mathrm{v}}+\hat{H}_{\mathrm{x}}$, is a sum of Hamiltonians $\hat{H}_{\mathrm{h,\alpha}}$ for each of the two horizontal legs $\alpha\in\{1,2\}$ of the ladder, the vertical coupling between them, $\hat{H}_{\mathrm{v}}$, and the cross-coupling $\hat{H}_{\mathrm{x}}$ present in experiment [39, 81]:
@@
$\hat{H}_{\mathrm{v}}$ $=$ $J\sum_{j}\left(\hat{\sigma}^{+}_{j,1}\hat{\sigma}^{-}_{j,2}+\mathrm{h.c.}\right),$ (1)
$\hat{H}_{\mathrm{x}}$ $=$ $\sum_{j}J_{\mathrm{x},j}\left(\hat{\sigma}^{+}_{j,1}\hat{\sigma}^{-}_{j+1,2}+\hat{\sigma}^{+}_{j,2}\hat{\sigma}^{-}_{j+1,1}+\mathrm{h.c.}\right),$
@@
where $\hat{\sigma}^{\pm}_{j,\alpha}$ is the standard raising (lowering) spin-$1/2$ operator for $j$th spin in layer $\alpha$. The ladder contains $L=2N$ spins in total. We use $\uparrow$, $\downarrow$’ to denote the standard spin basis states of a single site, assume open boundary conditions, and set $J=1$ as the overall energy scale. Note that the cross-coupling $J_{\mathrm{x},j}$ can be site-dependent or even random.
When $J_{\mathrm{x},j}\!=\!0$, the CL model hosts exact QMBS eigenstates which are simply the Dicke states of a collective spin of magnitude $N$:
$\hat{S}_{0}=\sum_{j=1}^{N}(|\!\!\Uparrow\rangle\langle\Downarrow\!\!|_{j}+|\!\!\Downarrow\rangle\langle\Uparrow\!\!|_{j}).$ (2)
[…]
Figure 1: (a)-(b) The Liouvillian spectra (left panel) and the overlap with $|\rho_{\Pi}\rangle\!\rangle$ (right panel) of the CL model, Eq. (Liouvillian Spectral Transition in Noisy Quantum Many-Body Scars), for dephasing strengths $\gamma\!=\!$ $0.0005$ (a) and $0.02$ (b). We plot the imaginary part of Liouvillian eigenvalues $\lambda_{k}$ against the real part of shifted eigenvalues $\lambda_{k}^{\prime}$. Red and blue dots represent scar (S) eigenvalues, while yellow and gray ones represent thermal (T) eigenvalues, labeled according to their inversion symmetry quantum number, $p\!=\!\pm 1$. Open black circles represent scar states corresponding to local maxima in the overlaps (right panels). The scar eigenvalues in $\mathrm{S},{-}1$’ sector undergo a spectral transition as $\gamma$ increases, while those in $\mathrm{S},{+}1$’ do not [compare red dots in panel (a) with (b)]. The data is obtained by exact diagonalization for couplings $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$, system size L=10L\!=\!10 with open boundary conditions.
Liouvillian spectral transition— We now consider each spin to have local decoherence to a bath. Under the Markov approximation, such a system is generally described by a Liouville superoperator [87]. Here, we focus on pure dephasing, with the corresponding superoperator
$\mathcal{L}=-i\left(\hat{H}\otimes\hat{I}-\hat{I}\otimes\hat{H}^{*}\right)+\gamma\mathcal{D},$ (3)
where $\gamma$ is the pure dephasing rate, the dissipator is $\mathcal{D}=\mathcal{D}^{\prime}-\sum_{j}\hat{I}_{j}\otimes\hat{I}_{j}$ with $\mathcal{D}^{\prime}=\sum_{j}\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}$. The operator $|E_{n}\rangle\langle E_{m}|$ in $\mathcal{L}$ has a vector representation given by $|E_{n}\rangle\!\otimes\!|E_{m}\rangle$ [87], where $|E_{n}\rangle$ and $|E_{m}\rangle$ are eigenstates of the entire Hamiltonian with $n,m=1,2,\dots,D$ and $D$ is the Hilbert space dimension. The $k$th eigenvalue and eigenmode of the Liouville superoperator in Eq. (3) are denoted by $\lambda_{k}$ and $|\lambda_{k}\rangle\!\rangle$, respectively, with $k=1,2,\cdots,D^{2}$. To understand the spectral properties of $\mathcal{L}$, we perform a transformation $\mathcal{L}^{\prime}\!=\!\mathcal{L}+\gamma\big(\mathrm{Tr}(\mathcal{D})/\mathrm{Tr}(\mathcal{I})\big)\mathcal{I}$, with $\mathcal{I}\!=\!\hat{I}\!\otimes\!\hat{I}$, which makes the Liouvillian spectrum traceless, i.e., sets the average of $\mathrm{Re}(\lambda^{\prime}_{k})$ to zero [82].
For an isolated system with cross-couplings turned off, $\gamma=J_{\mathrm{x},j}=0$, the Liouvillian of scar subspace is $\mathcal{S}_{0}=-i(\hat{S}_{0}\otimes\hat{I}-\hat{I}\otimes\hat{S}_{0})$. The spectrum of $\mathcal{S}_{0}$ is purely imaginary and the eigenmodes are tensor products of previously mentioned Dicke states for Eq. (2):
$\big|\lambda_{(l,s)}^{(0)}\big\rangle\!\big\rangle\!=\!\big|E_{l+s}^{\mathrm{S}}\big\rangle\!\otimes\!\big|E_{s}^{\mathrm{S}}\big\rangle,\quad\lambda_{(l,s)}^{(0)}\!=\!i\Big(E_{l+s}^{\mathrm{S}}\!-\!E_{s}^{\mathrm{S}}\Big)\!=\!2il\,$, (4)
where $E^{\mathrm{S}}$ and $|E^{\mathrm{S}}\rangle$ represent the eigenenergy and eigenstate of $\hat{S}_{0}$, respectively. For later convenience, we use indices $l\!=\!-N,-\!N{+}1,\cdots,N$ and $s\!=\!0,1,\cdots,N\!\!-\!|l|$ to label the scar eigenmodes. For small $\gamma$, we will use the overlap with states in Eq. (4) to identify QMBS eigenmodes of the full Liouvillian. For stronger dephasings, we identify QMBS states by following smooth changes of the spectrum as $\gamma$ is slowly varied.
The Liouvillian spectra of the CL model in Eq. (Liouvillian Spectral Transition in Noisy Quantum Many-Body Scars) with $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$ are presented in Fig. 1. We show illustrative examples of the spectra before (a) and after (b) the spectral transition. The QMBS eigenmodes (red and blue circles) are distinguished by their enhanced overlap with the state $|\rho_{\Pi}\rangle\!\rangle=|\Pi\rangle\!\otimes\!|\Pi\rangle$, also shown in Fig. 1. These eigenvalues are furthermore distinguished by their symmetry quantum number $p=\pm 1$ under inversion, generated by $\prod_{j}(\hat{\sigma}_{j,1}^{x}\hat{\sigma}_{j,2}^{x})\otimes(\hat{\sigma}_{j,1}^{x}\hat{\sigma}_{j,2}^{x})$. Before the spectral transition, Fig. 1(a), the identified QMBS eigenvalues cluster on the imaginary axis, unlike the thermal eigenvalues (yellow and gray circles). After the transition, Fig. 1(b), the same QMBS eigenvalues exhibit a qualitatively different distribution pattern in the form of a sparse diamond grid across the complex plane. For a system of $N$ qubits per leg, the QMBS eigenvalues undergoing this transition are layers $l=-N,2-N,\dots,N-2,N$ in Eq. (4). Thus, they alternate between inversion symmetry sectors given by $p=(-1)^{l}$.
The change in the distribution of eigenvalues in the complex plane in Fig. 1 is reminiscent of a LPTS breaking transition [82]. However, most of the thermal eigenvalues have already undergone LPTS breaking at $\gamma=0.0005$ in Fig. 1(a), suggesting that QMBS eigenvalues undergo a distinct transition. Moreover, it is only a subset of those eigenvalues belonging to $\mathrm{S},-1$’ sector in Fig. 1 that undergo such a transition. Below we show that the spectral transition of these QMBS eigenvalues leaves an imprint on the dissipative dynamics, and we develop their perturbative description to reveal the distinction from LPTS breaking transition of thermal eigenvalues.
@@
Liouvillian perturbation theory.—To examine the spectral transition in Fig. 1 more quantitatively, we study the spectra as we sweep the dephasing rate $\gamma$ in Fig. 2(a). At weak dephasing, several eigenvalues with inversion symmetry p=1p\!=\!-1 remain stationary along the real axis, while moving along the imaginary one. As $\gamma$ increases, pairs of these eigenvalues collide on the imaginary axis and abruptly bifurcate into two branches at specific dephasing rates $\gamma_{\star}^{\mathrm{S(1,2,3)}}$ along the real axis [for clarity, only the positive branch is shown in Fig. 2(a)].